# Monocular 3D Hand Pose Estimation with Implicit Camera Alignment

### v1

@@ Abstract
Estimating the 3D hand articulation from a single color image is a continuously investigated problem with applications in Augmented Reality (AR), Virtual Reality (VR), Human-Computer Interaction (HCI), and robotics. Apart from the absence of depth information, occlusions, articulation complexity, and the need for camera parameters knowledge pose additional challenges. In this work, we propose an optimization pipeline for estimating the 3D hand articulation from 2D keypoint input, which includes a keypoint alignment step and a fingertip loss to overcome the need to know or estimate the camera parameters. We evaluate our approach on the EgoDexter and Dexter+Object benchmarks to showcase that our approach performs competitively with the SotA, while also demonstrating its robustness when processing ”in-the-wild” images without any prior camera knowledge. Our quantitative analysis highlights the sensitivity of the 2D keypoint estimation accuracy, despite the use of hand priors. Code is available at https://github.com/cpantazop/HandRepo
@@ 1 Introduction
Reconstructing an articulated 3D hand from a single RGB image is a challenging problem in computer vision with a wide range of applications, including augmented and virtual reality (AR/VR), human-computer interaction (HCI), robotics, and the metaverse. However, it is a challenging task due to the lack of depth information, frequent object-related occlusions and self-occlusions, unknwon camera’s intrinsics parameters, and the hand’s complex articulation.
@@
    A camera-agnostic 3D hand pose estimation framework that leverages 2D keypoint detections,
@@
    A two-stage optimization pipeline combining rigid alignment and refinement with anatomical and fingertip constraints,
@@
    An extensive evaluation showcasing the method’s robustness in-the-wild and its competitive performance in standard benchmarks.
@@ 2 Prior Art
2D hand pose estimation. In 2D hand pose estimation, to estimate the pose of the hand we need to estimate the location of its keypoints. The human hand has 21 keypoints [10]: In Fig. 1 we see the keypoints on a hand. In prior work on this topic, in [27] Simon et al. present an approach that uses a multi-camera system to train fine-grained detectors for keypoints that are prone to occlusion, such as the joints of a hand. This procedure is called multiview bootstrapping. It uses an initial keypoint detector to generate noisy labels across multiple views, triangulates valid detections in 3D, and reprojects them as new training data to iteratively improve the detector. This process yields a real-time RGB keypoint detector with accuracy comparable to depth-based methods.
@@ 3 Background
where $r$ represents the residual error (i.e. , the difference between predicted and true values), and $\rho$ is a parameter that controls the sensitivity of the function to large residuals. It was originally introduced in the context of tomographic image reconstruction [13], and has since been applied in various domains, including computer vision and 3D human pose estimation [25].
Huber loss [14] combines MSE and MAE behavior and is defined as:
$L_{\delta}(a)=\begin{cases}\frac{1}{2}a^{2}&\text{for }|a|\leq\delta,\\ \delta\left(|a|-\frac{1}{2}\delta\right)&\text{otherwise},\end{cases}$ (4)
where $a=y-f(x)$. Let us now introduce a more specific loss constraint that is directly related to hand articulation and our problem. The anatomical joint limits error $E_{\text{limits}}$ is a penalty term applied to the 45-dimensional pose vector, where each joint has three DoF. This error function ensures that estimated hand poses remain within experimentally determined anatomical constraints, thereby enforcing plausible hand articulations. Similar to [11], this loss is implemented as a soft constraint using exponential functions that activate when joint angles exceed predefined limits:
@@
where $[l_{i},u_{i}]$ denote the lower and upper bounds for joint angle $\theta_{i}$, and $a_{\text{limits}}$ is an experimentally determined weight factor. The exponential terms enforce smooth constraints that discourage hand poses outside of the allowable range.
The concept of anatomical constraints for hand motion was first introduced by Lin  et al. [18] and has since been adopted by various works, including [11], [30], and others.
Rigid Transformations. A rigid transformation [6] involves applying a rotation, translation, and optionally scaling to align two sets of points while preserving their internal geometric relationships.
A rigid transformation consists of two main operations:
@@ 4 Method
Overview. To tackle the problem of fitting an articulated 3D hand from a single RGB image, we design the pipeline shown in Fig. 2. The input is a standard RGB image containing a human hand, and the output is a 3D hand in the exact same pose and orientation. The first step in the pipeline involves passing the input image through the MediaPipe Hand model estimator [1] to extract the 21 hand keypoints along with handedness information (whether the hand is left or right). This process results in a list of keypoints corresponding to pixel locations in the input image, which we later use as ground truth parameters for fitting the MANO hand model [26]. For the optimization step, our goal is to fit the MANO model’s keypoints to the extracted “ground truth” keypoints from MediaPipe. Given the pose and shape parameters of MANO, we can obtain the 21 hand joint locations through linear interpolation. MANO takes 45 pose parameters, 10 shape parameters, and 3 global rotation parameters as input to generate a 3D mesh that represents a unique hand configuration. The fitting process begins with a neutral “zero” pose and shape. Through a series of transformations and iterative optimization, we adjust the MANO parameters to align its 21 keypoints with the estimated MediaPipe keypoints by minimizing a loss function. A key challenge was when the input hand was in a different global rotation than MANO’s default pose. The root of the problem was in the initialization: MANO starts in a neutral pose, shape, and rotation since we have no prior information about the input hand’s orientation and initialized every parameter with zeros. However, if the hand in the input image had a significantly different rotation, for example, in a “handshake” position, the optimization process failed, resulting in completely implausible hand meshes.
Figure 3: Initialization and optimization pipeline.
Optimization Pipeline.
Table 1: The description of what combination of optimizer and loss function each experiment uses
@@
To address this, we compute a rigid transformation aligning the neutral MANO keypoints to MediaPipe keypoints. Additionally, we applied scaling to ensure both sets of keypoints were properly aligned before beginning the optimization process. The complete optimization pipeline is illustrated in Fig. 3. The rigid transformation was computed using six stable palm joints, namely keypoints [0,1,5,9,13,17] as illustrated in Fig. 1 to minimize the influence of finger articulation. To compute the transformation, we implemented a custom function that returns a $4\times 4$ transformation matrix, where the top-left $3\times 3$ block represents the rotation, the top-right $3\times 1$ column is the translation vector, and the bottom row is used for homogenous coordinates. For scaling, we computed a scale factor using the distance between keypoints 0 and 5 (wrist to index MetaCarpoPhalangeal joint) in both the target and MANO keypoints, ensuring anatomically proportional alignment. After the initial alignment, we used scipy.minimize [7] to optimize the MANO parameters, specifically leveraging either the BFGS, or the L-BFGS method. We experimented with three loss functions: MSE, GM, and Huber loss. To improve the accuracy of fingertip alignment, we explored weighted loss functions, as we observed that most keypoints were densely concentrated around the palm. Due to the nature of the loss functions, the optimization process primarily focused on minimizing the error in these denser regions, often leading to less precise alignment of the fingertips. However, in real-world hand movement, fingertips play a critical role in defining hand gestures and poses, making their accurate positioning essential. To address this, we applied weighted loss functions across all three variations (MSE, GM, and Huber), giving higher importance to fingertip keypoints to ensure their proper alignment. Additionally, we attempted to integrate anatomical joint constraints into the loss function to enforce physically plausible hand poses. However, this approach proved ineffective, as it restricted the optimization process too severely. Despite experimenting with various weighting schemes, the results did not improve. Instead, we adopted a two-stage optimization strategy:
@@
After optimization, we reversed the rigid transformation to recover the MANO keypoints and mesh in the original scale, orientation, and position. The rotation matrix, translation vector, and scaling factor used for initialization were inverted to map the optimized MANO results back to the target’s coordinate system so that we can inspect our results visually. To determine the root pose, we computed the axis-angle representation of the rotation matrix derived from the rigid transformation and incorporated it into the MANO model’s parameters. This ensured that the global rotation was accurately represented in the final output. This means that we now also have a 3D hand with the correct global orientation and pose but with the zero shape that is not really a problem.
@@ 5.1 Ablation Study
Tab. 1 presents the different experimental configurations we tested to evaluate our method on two datasets. Each experiment is identified by a letter (ID) and corresponds to a specific combination of optimizers and loss functions. While we tested all loss function combinations using the L-BFGS optimizer, we selectively tested what we considered the most promising loss function setup with the BFGS optimizer. Tab. 2 presents the quantitative results of our experiments, reporting the End-Point Error (in mm) and the AUC of PCK for both datasets. Our findings indicate that most experiments perform consistently well, with results comparable to each other. This consistency is expected, as the different loss functions are all variations of the MSE loss, designed to improve robustness against outliers. The best-performing results for each metric are highlighted in bold, although in many cases, multiple configurations achieve similar results. Our observations suggest that weighted fingertips loss functions tend to yield slightly better performance across metrics. However, the standard loss functions also achieve competitive results, while the better results of the weighted fingertips loss may be attributed to the specific evaluation datasets focusing primarily on fingertip keypoints. Additionally, we can observe that in experiment H, this two-stage approach integrating anatomical constraints resulted in worse overall metrics although it successfully corrected depth-related errors caused by MediaPipe’s 3D predictions in specific cases.
Overall, our results suggest that the most effective optimizer-loss combinations were Experiment A and Experiment G. This outcome aligns with expectations, as L-BFGS and BFGS are closely related, with L-BFGS being a memory-efficient variant. If computational efficiency is considered alongside accuracy, Experiment A—using L-BFGS with a weighted combination of MSE losses—emerges as the optimal choice. Figs. 4(a) and 4(b) compare the performance of the methods discussed above on the EgoDexter and Dexter+Object datasets, respectively. In both cases, the method incorporating anatomical constraints performs the worst. For the EgoDexter dataset, we observe that methods utilizing weighted fingertip loss functions consistently outperform those with standard loss functions, forming a distinct cluster of higher-performing models. However, in the Dexter+Object dataset, this distinction is less pronounced.
(a) 3D PCK for EgoDexter dataset.
(b) 3D PCK for Dexter+Object dataset.
@@ 5.2 Quantitative Evaluation
Table 3: AUC of PCK ($\uparrow$) comparison with state-of-the-art methods. We use “*” to note the methods that work in real-time that are more challenging.
@@ 5.3 Qualitative Evaluation
Finally, in Fig. 5(f), we demonstrate that our method generalizes beyond structured datasets by estimating hand poses from an image of the Mona Lisa painting. This proves that our approach does not rely on camera parameters and can function on in-the-wild RGB images, making it applicable in diverse real-world scenarios.
@@ 6 Conclusion
In this paper, we proposed an optimization-based solution for estimating the 3D articulation of a human hand from a single RGB image, without knowning the camera intrinsic parameters. Our method leveraged the MediaPipe keypoint detector to obtain an initial estimation of the hand joints in 2D space, and performed a fitting stage using the MANO parametric model to obtain the 3D joint rotations. For the fitting stage we incorporated a fingertip alignment loss coupled with anatomical constraints. Our extensive evaluation demonstrated that our approach can robustly operate in-the-wild without the need for prior camera parameter information, while being competitive when compared to SotA data-driven models.

### v2

@@ Abstract
Estimating the 3D hand articulation from a single color image is an important problem with applications in Augmented Reality (AR), Virtual Reality (VR), Human-Computer Interaction (HCI), and robotics. Apart from the absence of depth information, occlusions, articulation complexity, and the need for camera parameters knowledge pose additional challenges. In this work, we propose an optimization pipeline for estimating the 3D hand articulation from 2D keypoint input, which includes a keypoint alignment step and a fingertip loss to overcome the need to know or estimate the camera parameters. We evaluate our approach on the EgoDexter and Dexter+Object benchmarks to showcase that it performs competitively with the state-of-the-art, while also demonstrating its robustness when processing “in-the-wild” images without any prior camera knowledge. Our quantitative analysis highlights the sensitivity of the 2D keypoint estimation accuracy, despite the use of hand priors. Code is available at the project page https://cpantazop.github.io/HandRepo/
@@ 1 Introduction
Reconstructing an articulated 3D hand from a single RGB image is a challenging problem in computer vision with a wide range of applications, including augmented and virtual reality (AR/VR), human-computer interaction (HCI), robotics, and the metaverse. However, it is a challenging task due to the lack of depth information, frequent object-related occlusions and self-occlusions, unknown camera intrinsics parameters, and the hand’s complex articulation.
@@
    A camera-agnostic 3D hand pose estimation framework that leverages 2D keypoint detections.
@@
    A two-stage optimization pipeline combining rigid alignment and refinement with anatomical and fingertip constraints.
@@
    An extensive evaluation showcasing the method’s robustness in-the-wild and its competitive performance on standard benchmarks.
@@ 2 Prior Art
2D hand pose estimation. In 2D hand pose estimation, to estimate the pose of the hand we need to estimate the location of its keypoints. The human hand has 21 keypoints [10]: In Fig. 1 we depict the keypoints on a hand. In prior work on this topic, in [27] Simon et al. present an approach that uses a multi-camera system to train fine-grained detectors for keypoints that are prone to occlusion, such as the joints of a hand. This procedure is called multiview bootstrapping. It uses an initial keypoint detector to generate noisy labels across multiple views, triangulates valid detections in 3D, and reprojects them as new training data to iteratively improve the detector. This process yields a real-time RGB keypoint detector with accuracy comparable to depth-based methods.
@@ 3 Background
where $r$ represents the residual error (i.e., the difference between predicted and true values), and $\rho$ is a parameter that controls the sensitivity of the function to large residuals. It was originally introduced in the context of tomographic image reconstruction [13], and has since been applied in various domains, including the 3D human pose estimation [25].
The Huber loss [14] combines the MSE and MAE behavior and is defined as:
$L_{\delta}(a)=\begin{cases}\frac{1}{2}\>\ a^{2}&\text{for }|a|\leq\delta,\\ \delta\left(|a|-\frac{1}{2}\delta\right)&\text{otherwise},\end{cases}$ (4)
where $a=y-f(x)$.
Let us now introduce a more specific loss constraint that is directly related to hand articulation and our problem. The anatomical joint limits error $E_{\text{limits}}$ is a penalty term applied to the 45-dimensional pose vector, where each joint has three DoF. This error function ensures that estimated hand poses remain within experimentally determined anatomical constraints, thereby enforcing plausible hand articulations. Similar to [11], this loss is implemented as a soft constraint using exponential functions that activate when joint angles exceed predefined limits:
@@
where $[l_{i},u_{i}]$ denote the lower and upper bounds for joint angle $\theta_{i}$, and $a_{\text{limits}}$ is an experimentally determined weight factor. The exponential terms enforce smooth constraints that discourage hand poses outside the allowable range. The concept of anatomical constraints for hand motion was first introduced by Lin  et al. [18] and has since been adopted by various works, including [11], [30], and others.
Rigid Transformations. A rigid transformation [6] involves applying a rotation, translation, and optionally scaling to align two sets of points while preserving their internal geometric relationships. A rigid transformation consists of two main operations:
@@ 4 Method
Overview. To tackle the problem of fitting an articulated 3D hand from a single RGB image, we design the pipeline shown in Fig. 2. The input is a standard RGB image containing a human hand, and the output is a 3D hand in the exact same pose and orientation. The first step in the pipeline involves passing the input image through the MediaPipe Hand model estimator [1] to extract the 21 hand keypoints along with handedness information (whether the hand is left or right). This process results in a list of keypoints corresponding to pixel locations in the input image, which we later use as ground truth parameters for fitting the MANO hand model [26]. For the optimization step, our goal is to fit the MANO model’s keypoints to the extracted “ground truth” keypoints from MediaPipe. Given the pose and shape parameters of MANO, we can obtain the 21 hand joint locations through linear interpolation. MANO takes 45 pose parameters, 10 shape parameters, and 3 global rotation parameters as input to generate a 3D mesh that represents a unique hand configuration. The fitting process begins with a neutral “zero” pose and shape. Through a series of transformations and iterative optimization, we adjust the MANO parameters to align its 21 keypoints with the estimated MediaPipe keypoints by minimizing a loss function. A key challenge was when the input hand was in a different global rotation than MANO’s default pose. The root of the problem was the initialization: MANO starts in a neutral pose, shape, and rotation, since we have no prior information about the input hand’s orientation, and initialized every parameter with zeros. However, if the hand in the input image had a significantly different rotation, for example, in a “handshake” position, the optimization process failed, resulting in completely implausible hand meshes.
Figure 3: Optimization pipeline (MP stands for MediaPipe).
Table 1: The description of what combination of optimizer and loss function each experiment uses.
@@
Optimization Pipeline. To address this challenge, we compute a rigid transformation aligning the neutral MANO keypoints to MediaPipe keypoints. Additionally, we applied scaling to ensure both sets of keypoints were properly aligned before beginning the optimization process. The complete optimization pipeline is illustrated in Fig. 3. The rigid transformation was computed using six stable palm joints, namely keypoints [0,1,5,9,13,17] as illustrated in Fig. 1 to minimize the influence of finger articulation. To compute the transformation, we implemented a custom function that returns a $4\times 4$ transformation matrix, where the top-left $3\times 3$ block represents the rotation, the top-right $3\times 1$ column is the translation vector, and the bottom row is used for homogenous coordinates. For scaling, we computed a scale factor using the distance between keypoints 0 and 5 (wrist to index MetaCarpoPhalangeal joint) in both the target and MANO keypoints, ensuring anatomically proportional alignment. After the initial alignment, we used scipy.minimize [7] to optimize the MANO parameters, specifically leveraging either the BFGS, or the L-BFGS method. We experimented with three loss functions: MSE, GM, and Huber loss. To improve the accuracy of fingertip alignment, we explored weighted loss functions, as we observed that most keypoints were densely concentrated around the palm. Due to the nature of the loss functions, the optimization process primarily focused on minimizing the error in these denser regions, often leading to less precise alignment of the fingertips. However, in real-world hand movement, fingertips play a critical role in defining hand gestures and poses, making their accurate positioning essential. To address this, we applied weighted loss functions across all three variations (MSE, GM, and Huber), giving higher importance to fingertip keypoints to ensure their proper alignment. Additionally, we attempted to integrate anatomical joint constraints into the loss function to enforce physically plausible hand poses. However, this approach proved ineffective, as it restricted the optimization process too severely. Despite experimenting with various weighting schemes, the results did not improve. Instead, we adopted a two-stage optimization strategy:
@@
After optimization, we reversed the rigid transformation to recover the MANO keypoints and mesh in the original scale, orientation, and position. The rotation matrix, translation vector, and scaling factor used for initialization were inverted to map the optimized MANO results back to the target’s coordinate system so that we can inspect our results visually. To determine the root pose, we computed the axis-angle representation of the rotation matrix derived from the rigid transformation and incorporated it into the MANO model parameters. This ensured that the global rotation was accurately represented in the final output. This means that we now also have a 3D hand with the correct global orientation and pose but with the zero shape that is not really a problem.
@@ 5.1 Ablation Study
Tab. 1 presents the different experimental configurations we tested to evaluate our method on two datasets. Each experiment is identified by a letter (ID) and corresponds to a specific combination of optimizers and loss functions. While we tested all loss function combinations using the L-BFGS optimizer, we selectively tested what we considered the most promising loss function setup with the BFGS optimizer. Tab. 2 presents the quantitative results of our experiments, reporting the End-Point Error (in mm) and the AUC of PCK for both datasets. Our findings indicate that most experiments perform consistently well, with results comparable to each other. This consistency is expected, as the different loss functions are all variations of the MSE loss, designed to improve robustness against outliers. The best-performing results for each metric are highlighted in bold, although in many cases, multiple configurations achieve similar results. Our observations suggest that weighted fingertips loss functions tend to yield slightly better performance across metrics. However, the standard loss functions also achieve competitive results, while the better results of the weighted fingertips loss may be attributed to the specific evaluation datasets focusing primarily on fingertip keypoints. Additionally, we can observe that in experiment H, this two-stage approach integrating anatomical constraints resulted in worse overall performance although it successfully corrected depth-related errors caused by MediaPipe’s 3D predictions in specific cases.
Overall, our results suggest that the most effective optimizer-loss combinations were Experiment A and Experiment G. This outcome aligns with expectations, as L-BFGS and BFGS are closely related, with L-BFGS being a memory-efficient variant. If computational efficiency is considered alongside accuracy, Experiment A, using L-BFGS with a weighted combination of MSE losses, emerges as the optimal choice. Figs. 4(a) and 4(b) compare the performance of the methods discussed above on the EgoDexter and Dexter+Object datasets, respectively. In both cases, the method incorporating anatomical constraints performs the worst. For the EgoDexter dataset, we observe that methods utilizing weighted fingertip loss functions consistently outperform those with standard loss functions, forming a distinct cluster of higher-performing models. However, on the Dexter+Object dataset, this distinction is less pronounced.
(a) 3D PCK results on the EgoDexter dataset.
(b) 3D PCK results on the Dexter+Object dataset.
@@ 5.2 Quantitative Evaluation
Table 3: AUC of PCK ($\uparrow$) comparison with state-of-the-art methods. We use “*” to note the methods that work in real-time, which is a more challenging task.
@@ 5.3 Qualitative Evaluation
Finally, in Fig. 5(f), we demonstrate that our method generalizes beyond structured datasets by estimating hand poses from an image of the Mona Lisa painting. This showcases that our approach does not rely on camera parameters and can function on in-the-wild RGB images, making it applicable in diverse real-world scenarios.
@@ 6 Conclusion
In this paper, we proposed an optimization-based solution for estimating the 3D articulation of a human hand from a single RGB image, without knowledge of the camera intrinsic parameters. Our method leveraged the MediaPipe keypoint detector to obtain an initial estimation of the hand joints in the 2D space, and it performed a fitting stage using the MANO parametric model to obtain the 3D joint rotations. For the fitting stage we incorporated a fingertip alignment loss coupled with anatomical constraints. Our extensive evaluation demonstrated that our approach can robustly operate in-the-wild without the need for prior camera parameter information, while being competitive when compared to SotA data-driven models.